### v1

# Moral Alignment for LLM Agents

Abstract
Decision-making agents based on pre-trained Large Language Models (LLMs) are increasingly being deployed across various domains of human activity. While their applications are currently rather specialized, several research efforts are under way to develop more generalist agents. As LLM-based systems become more agentic, their influence on human activity will grow and the transparency of this will decrease. Consequently, developing effective methods for aligning them to human values is vital.
The prevailing practice in alignment often relies on human preference data (e.g., in RLHF or DPO), in which values are implicit and are essentially deduced from relative preferences over different model outputs. In this work, instead of relying on human feedback, we introduce the design of reward functions that explicitly encode core human values for Reinforcement Learning-based fine-tuning of foundation agent models. Specifically, we use intrinsic rewards for the moral alignment of LLM agents.
We evaluate our approach using the traditional philosophical frameworks of Deontological Ethics and Utilitarianism, quantifying moral rewards for agents in terms of actions and consequences on the Iterated Prisoner’s Dilemma (IPD) environment. We also show how moral fine-tuning can be deployed to enable an agent to unlearn a previously developed selfish strategy. Finally, we find that certain moral strategies learned on the IPD game generalize to several other matrix game environments. In summary, we demonstrate that fine-tuning with intrinsic rewards is a promising general solution for aligning LLM agents to human values, and it might represent a more transparent and cost-effective alternative to currently predominant alignment techniques.
1 Introduction
The alignment problem is an active field of research in Machine Learning (Christian, 2020; Weidinger et al., 2021; Anwar et al., 2024; Gabriel et al., 2024; Ji et al., 2024; Ngo et al., 2024). It is gaining even wider importance with the advances and rapid deployment of Large Language Models (LLMs, Anthropic 2024; Gemini Team 2024; OpenAI 2024). The most common practices in the alignment of LLMs today involve Reinforcement Learning from Human Feedback (RLHF - Glaese et al. 2022; Ouyang et al. 2022; Bai et al. 2023) or Direct Preference Optimization (DPO - Rafailov et al. 2024). Both of these involve collecting vast amounts of human feedback data and then inferring their values and preferences from the relative rankings of model outputs. In doing so, human values are implicitly represented.
This approach poses certain challenges (Casper et al., 2023). Specifically, collecting preference data is very costly and often relies on potentially unrepresentative samples of human raters. Indeed, the values derived through this process are strongly dependent on the selection criteria of the pool of individuals. Furthermore, human preferences are notoriously complex and inconsistent. In RLHF, the values that are ultimately incorporated into the fine-tuned models are learned by a reward model from data in a fully bottom-up fashion, and are never made explicit to any human oversight. One might argue that current LLMs fine-tuned with these methods are able to provide “honest, harmless and helpful” responses (Glaese et al., 2022; Bai et al., 2023) and already display certain moral values (Schramowski et al., 2022; Abdulhai et al., 2023; Hartmann et al., 2023). Methods to guide the social behavior of LLM agents via interaction in rule-guided simulated societies have also been proposed recently (Liu et al., 2024). However, models’ apparent values can also be interpreted as “moral mimicry” of their users when responding to these prompts (Simmons, 2022; Shanahan et al., 2023). As a consequence, given phenomena such as situationally-aware reward-hacking or misalignment in internally-represented goals (Ngo et al., 2024), the true values learned by the models through these methods may give rise to dangerous behaviors, which will not be explicitly known until after deployment.
Our work aims to address this type of goal misgeneralization in particular by providing clearer, explicit moral alignment goals as intrinsic rewards for fine-tuning RL-based algorithms [^1] For a more comprehensive discussion of learning as a method for moral alignment with implicit (bottom-up) versus explicit (top-down) principles, we refer the interested reader to Tennant et al. (2023b).. In this study, we approach alignment from an agent-based perspective. Since LLMs are increasingly adopted as a basis for strategic decision-making systems and agentic workflows (Wang et al., 2024b), it is critical that we align the choices made by LLM agents with our values, including value judgments about what actions are morally good or bad (Amodei et al., 2016; Anwar et al., 2024). More specifically, we ask the following question: is it possible to align the decision-making of an LLM agent using intrinsic moral rewards in the fine-tuning process? Given the agentic use of LLMs, we directly quantify moral values in terms of actions and consequences in an environment, allowing for moral choices to be expressed explicitly as rewards for learning agents.
We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” a selfish strategy. If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values. A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain. At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others. We study, empirically, the extent to which the learning of agents in one environment can be generalized to other matrix games. In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications.
To summarize, our study provides the following contributions:
  • We introduce a novel, general solution for aligning LLM agents to human moral values by means of fine-tuning via Reinforcement Learning with Intrinsic Rewards.
  • We evaluate the approach using a repeated social dilemma game environment (with fixed-strategy and learning opponents), and Deontological and Utilitarian moral values. We show that LLM agents fine-tuned with intrinsic rewards are able to successfully learn aligned moral strategies.
  • We discuss how the proposed solution can be generalized and applied to different scenarios in which moral choices can be captured by means of payoff matrices.
2 Background
2.1 LLM Agents
Agency refers to the ability of a system to decide to take actions in the world (Swanepoel & Corks, 2024). In this paper, we equate agency with strategic decision-making - i.e., making a choice in an environment in which multiple actions are available and lead to different outcomes. In the case of LLMs, this view assumes that model outputs will be interpreted as actions in some environment. The simplest way of implementing this is through the use of specific tokens to represent the actions. Particular tokens can be reserved or fine-tuned from the model’s vocabulary to represent actions, and planning and reasoning ability can be improved via action-driven prompting strategies (Yao et al., 2023). Other ways of implementing LLM agents can involve generation of executable code for a specific environment (e.g., a video game, Wang et al. 2024a) or connection to various tool APIs (e.g., Patil et al. 2023; Shen et al. 2023), but these are more specialized and, therefore, not the focus of this work.
Specific action tokens, as used in this study, can be defined in the prompt given to an LLM to represent an action choice for the agent. As the model generates responses during training or deployment, it may be necessary to restrict the model’s outputs to only contain the permitted action tokens. Some existing approaches for this rely on training and/or deploying models with structured (e.g., JSON) output formats or constrained generation (Beurer-Kellner et al., 2024), which suppresses the probabilities of all tokens in the model’s output layer except for the legal action tokens. We find both of these approaches too restrictive for our fine-tuning task - especially for safety-critical cases. Fine-tuning based on a restricted output space or format poses risks of the model “hiding” undesirable behaviors (Anwar et al., 2024). Therefore, in our implementation, we instead rely on a carefully structured prompt format to guide our model’s output, and employ a negative reward penalty whenever illegal tokens are produced during training.
Using the techniques outlined, agents based on pre-trained LLMs combined with other elements of various cognitive architectures (Sumers et al., 2024) such as a skill set (Wang et al., 2024a) or a memory store (Vezhnevets et al., 2023) have been used to reasonably simulate decision-making in open-ended environments (Wang et al., 2024b), including those with only a single-agent (Wang et al., 2024a) or of a multi-agent nature (Park et al., 2023). Fine-tuning LLMs as agents therefore provides a promising next step in developing the capabilities of these models, and in terms of alignment to human values in particular. LLMs fine-tuned with RLHF, and especially those fine-tuned to follow human instructions, have been shown to become more goal-directed than simple sequence-completion foundation models (Glaese et al., 2022; Ouyang et al., 2022; Bai et al., 2023). We rely on instruction-tuned LLMs in this study and use the Gemma2-2b-it model in particular (Gemma Team, 2024) as a decision-making agent in social dilemma games. Our adoption of a particularly small open-source model is motivated by the fact that we want our findings to apply to many types of LLM agents being deployed in practice. Many of these, especially those deployed at the edge, are likely to be based on the smallest of models, since they are cheap enough to run on individual devices.
2.2 Fine-tuning LLM Agents with Reinforcement Learning
Proximal Policy Optimization (PPO, Schulman et al. 2017) is the most commonly used technique for fine-tuning LLMs with RL (Stiennon et al., 2022). This on-policy method is often deployed in conjunction with a Kullback-Leibler (KL) penalty to prevent the new model from shifting too far away from the original underlying token distribution and thus losing other capabilities such as producing coherent linguistic output (Jaques et al., 2017; Ziegler et al., 2020; Stiennon et al., 2022). Offline fine-tuning methods have also been developed (Snell et al., 2023) and may provide a more sample-efficient alternative in the future. The reward signal for RL-based training in existing implementations tends to be derived from preference data provided by human raters (Glaese et al., 2022; Ouyang et al., 2022; Bai et al., 2023) or a constitution of other human and/or artificial agents (Bai et al., 2022; Huang et al., 2024). In this study we propose a new methodology for RL-based fine-tuning with intrinsic moral rewards.
Compared to non-linguistic RL agent training, the pre-trained LLM in this case can be viewed as providing a common-sense model  [^2] We note that the extent of true commonsense knowledge of LLMs is still debated (Mitchell, 2021), especially for smaller models. Nevertheless, benchmark evaluations suggest the emergence of common sense and reasoning abilities even in models as small as 2b parameters - for example, Gemma2-2b-it scores over 85% (Gemma Team, 2024) on the commonsense benchmark introduced by Zellers et al. 2019. of the world (Wong et al., 2023), equipping an LLM-based agent with some intuition about potential dynamics of various environments. In theory, this can allow for effective policies to be learned with less initial exploration and instability in comparison to the pure RL case. Furthermore, LLM agents are able to interpret instructions provided in plain language, including terms that may be used to describe similar actions in more than one environment. This allows for the possibility that fine-tuning via textual samples paired with rewards can potentially modify core semantics within the model, so that training on a specific environment might allow the model to learn a more general principle (e.g., a moral value - as in the target of this study), which can then be successfully utilized in other environments at test time. Early results from text-instructed video models suggest that this generalization of learned behaviors across environments is indeed possible (SIMA Team, 2024). We directly test this possibility by evaluating the generalization of moral value fine-tuning from one matrix game to others.
2.3 Social Dilemma Games
A prominent social dilemma game is the Iterated Prisoner’s Dilemma (IPD), in which a player can Cooperate (C) with their opponent for mutual benefit, or betray them - i.e., Defect (D) for individual reward (Rapoport, 1974; Axelrod & Hamilton, 1981). The payoffs in any step of the game are determined by a payoff matrix, presented for the row player versus a column player in Figure 1.
C D
C 3,3 0,4
D 4,0 1,1
Figure 1: Payoffs for the Iterated Prisoner’s Dilemma
In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C). Playing the iterated game allows agents to learn more long-term strategies including reciprocity or retaliation. While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas). This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
The behavior of LLM agents on decision-making and game-theoretic scenarios has been extensively debated in recent literature (Gandhi et al., 2023; Fan et al., 2024; Zhang et al., 2024). LLM agents have been found to act differently to humans, and in ways that are still not fully “rational” in terms of forming goals from a prompt, refining beliefs, or taking optimal actions based on those goals and beliefs (Fan et al., 2024; Macmillan-Scott & Musolesi, 2024). Large-scale state-of-the-art models playing the IPD have been observed to deploy sensible yet “unforgiving” strategies (Akata et al., 2023), but some benchmark datasets suggest that these models lack true strategic reasoning in games including the IPD (Duan et al., 2024). New developments in in-token reasoning capabilities of state-of-the-art LLM-based platforms (OpenAI, 2024) as well as prompting strategies specifically centered around reasoning and acting (Wei et al., 2022; Shinn et al., 2023; Yao et al., 2023) are likely to improve these capabilities, though existing results suggest that the benefits of these methods are more likely to arise for very large foundation models (Bubeck et al., 2023). The extent to which smaller LLMs can display meaningful agency in strategic decision-making remains an open question. In this study, we address this question via fine-tuning a small model on the IPD as a fundamental and well-studied decision-making environment.
2.4 Intrinsic Rewards for Moral Alignment
In this work, we directly specify alignment goals for agents by defining intrinsic rewards in terms of actions in a social dilemma environment. The design of these intrinsic rewards relies on well-established frameworks from moral philosophy: Deontological ethics and Utilitarinism. Deontological ethics (Kant, 1785) considers an agent moral if their actions conform to certain norms. A prominent example of a norm is conditional cooperation (i.e., “it is unethical to defect against a cooperator"). This norm forms an essential component of direct and indirect reciprocity, a potentially essential mechanism for the evolution of cooperation in human and animal societies (Nowak, 2006). Utilitarian morality (Bentham, 1780), on the other hand, is a type of consequentialist reasoning, according to which an agent is deemed moral if their actions maximize collective “welfare” for all agents in their society (or, in this case, collective payoff for all players in the game), and less attention is paid to whether current actions adhere to norms. Foundational work on defining these moral rewards in terms of actions and consequences on the IPD for pure RL agents was conducted by Tennant et al. (2023a) and Tennant et al. (2024). In this paper, we evaluate the extent to which this framework can be applied to align the behavior of LLM-based ones.
3 Fine-tuning Methodology
3.1 Agent and Environment
The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game. At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move. Within the MDP framework, each player’s current action affects the game’s state at the next time step.
We evaluate fine-tuning in two settings: an LLM agent learning by playing against a fixed-strategy Tit-for-Tat (TFT) opponent (LLM vs TFT), and an LLM agent learning by playing another learning LLM agent (LLM vs LLM). We chose TFT as a specific type of fixed-strategy opponent from the literature given its characteristics, i.e., being forgiving, defensive and, at the same time, interpretable (Axelrod & Hamilton, 1981; Binmore, 2005). Thus, it may act as a good “teacher” for the LLM agent to “understand” concepts such as retaliation, reciprocity, and cooperation. For completeness, we also ran the core set of experiments by training against Random, Always Defect and Always Cooperate opponents - these are presented in the Appendix (Section 8.5). The LLM vs LLM case is a more complex scenario that may lead to non-stationarity due to two separate models being updated continuously, but which also presents great interest due to the difficulty in predicting the outcomes from multi-agent learning (Busoniu et al., 2008).
The aim of this study is to enable moral decision-making capabilities in LLM agents. We perform fine-tuning based on a single environment - the IPD. However, we aim to mobilize the general decision-making elements of the model in playing the game, rather than allowing it to retrieve memorized responses for the Prisoner’s Dilemma that were present in its pre-training data. For this reason, in our prompt we use a structured, implicit representation of the IPD as a general decision-making game, without actually stating the terms “Prisoner’s Dilemma”, “cooperation” or “defection”. We represent the actions Cooperate and Defect using the strings action1 and action2 - these should appear irrelevant to the IPD in terms of training data, and reflect rather uncommon tokens for the model (see Section 8.2 in the Appendix for an illustration of the prompt). Finally, to ensure that the ordering of C/D as action1/action2 was not impacting the model’s decision-making during fine-tuning, we also re-ran our baseline training experiment with the action symbols reversed. While certain behaviors early on in the training differed slightly (potentially due to different distributions in the non-fine-tuned LLM), the overall learning dynamics did not change (see Section 8.4 in the Appendix for the results).
3.2 Moral Fine-tuning Procedure
Table 1: Definitions of the types of moral rewards used in fine-tuning the LLM agent, from the point of view of the moral agent $M$ playing versus an opponent $O$ at time step $t$.
Moral Fine-tuning Type Moral Reward Function
Game reward (selfish) $R_{M}^{t}=\begin{cases}R^{t}_{M_{\text{game}}},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Deontological reward $R_{M}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Utilitarian reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}+R_{O_{\text{game}}}^{t},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
Game+Deontological reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ R_{M_{\text{game}}}^{t},&\text{otherwise if}a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
We run training in $T$ episodes: each episode begins with a random state being incorporated into the IPD prompt. The LLM-based agent $M$ then plays $N$ repetitions of the IPD game against an opponent $O$ (where $N$ is the batch size). On each repetition, the players’ actions from the previous time step are reflected in their opponent’s current state $s^{t}_{M}=(a^{t-1}_{O},a^{t-1}_{M})$. If an LLM player outputs an illegal move on a certain time step, this move is not used to update their opponent’s state, but the agent still learns from the experience. After $N$ games have been played, the LLM agent performs a PPO learning step update based on the gathered batch of experiences. This marks the end of an episode.
In our core experiments, we test four different reward signals for moral fine-tuning of LLM agents (as outlined in Table 1): 1) the Game reward $R^{t}_{M_{\text{game}}}$, representing the goals of a selfish or rational agent playing the IPD 2) a Deontological reward $-\xi$ for violating the moral norm “do not defect against an opponent who previously cooperated”, 3) Utilitarian reward, representing the collective payoff in the game, and 4) Game+Deontological reward which combine game payoff with a Deontological penalty in a multi-objective manner. In addition, we test whether a model fine-tuned on Game rewards is able to unlearn this selfish strategy via further fine-tuning with moral rewards. Therefore, we additionally fine-tune agents with: 5) Game, then Deontological reward (training with each reward type for half of the total number of episodes $T$), and 6) Game, then Utilitarian reward (again, each for half of the total duration of $T$ episodes). Finally, during each type of fine-tuning we also implement a penalty $R_{\text{illegal}}$ for generating “illegal” action tokens, to encourage the model to keep its answers within the permitted action space, as defined in the game prompt.
3.3 Implementation Details
We use Gemma2-2b-it (Gemma Team, 2024) as our core agent model to be fine-tuned, being one of the most popular and performant small open-source models. The small model footprint allows us to run computationally feasible experiments through the use of LoRA (Hu et al., 2022) and 4-bit quantization. We use the TRL library (von Werra et al., 2020) to fine-tune the LLM with PPO. We run training for $T=1000$ episodes for each fine-tuning variation. In our PPO implementation, we use batch sizes of $N=3$ and $N=5$ for LLM vs LLM and LLM vs TFT training, respectively, which strikes a nice balance between not running out of available CUDA memory, yet providing sufficient experience for stable and efficient training  [^3] The code implementing the model fine-tuning and analysis will be made available upon acceptance. We use reward scaling and normalization (Engstrom et al., 2020), as well as gradient accumulation with 4 steps, and employ LoRA (Hu et al., 2022) with rank 64, so that the total number of parameters we train is only around 5% of the original model. Otherwise, we keep all PPO parameters as their default values in the TRL package, including the optimizer’s learning rate and adaptive KL control (Jaques et al., 2017). In terms of reward parameters, we set $\xi=3$ and $R_{\text{illegal}}=-6$. We select the tokens action1 and action2 as the only “legal” tokens in response to the IPD prompt ($C_{\text{legal}}=$ action1, $D_{\text{legal}}=$ action2). The action symbols are each encoded as two tokens in the model’s tokenizer, so during training we restrict the maximum output length for model generations to two tokens. For detail on parameter selection, please refer to Appendix 8.1.
4 Evaluating the effectiveness of fine-tuning: moral choices on the IPD
4.1 Evaluation Approach
First of all, we analyze the learning dynamics observed as models develop the ability to meet the moral goals set in their rewards (Section 4.2). We analyze learning against a static opponent and a learning opponent. We then assess the effectiveness of moral “unlearning” (Section 4.3). Beyond measuring success on IPD fine-tuning itself, we evaluate the generalization of the moral fine-tuning from one matrix game environment onto four other matrix games (Section 5.1): Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game. The payoffs and further details around these games are presented in the Appendix (Section 8.6). Finally, we evaluate the extent to which fine-tuning on the IPD alters the models’ behavior on more general prompts, as well as the explicit IPD game (Section 5.2). For each experiment, we report average results across five random seeds.
4.2 Learning dynamics
a) Refer to caption Refer to caption Refer to caption Refer to caption
b) Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
In general, we find that it is possible to fine-tune the LLM agents to choose actions that are consistent with certain moral and/or game rewards in the IPD. We analyze learning dynamics over the four core types of fine-tuning in Figure 2. Learning against a fixed-strategy opponent (panel a), fine-tuning on Game rewards (i.e., rewards assigned through the payoff matrix of the game), the agent learns a defective policy, which forms a Nash Equilibrium versus a TFT opponent (Axelrod & Hamilton, 1981). In the case of Deontological fine-tuning, the agent quickly learns to avoid defecting against a cooperator nearly 100% of the time, thus never violating the moral norm encoded in the respective reward function. In practice, this agent also learns to prefer cooperation in general, though this was not directly encouraged by the Deontological norm (in terms of Deontological reward, defecting against a defector is just as valid as cooperating against a cooperator - see reward definition in Table 1). On Utilitarian fine-tuning, the agent learns to achieve mutual cooperation against a TFT opponent, which is expected given that this strategy offers the optimal way to obtain the highest collective reward on the IPD. Finally, in the case of fine-tuning with a multi-objective Game+Deontological reward, the agent learns a 50%-50% Cooperate-Defect strategy, but also learns to avoid defecting against a cooperator. Thus, this agent does not violate their moral norm even as they work to obtain high payoffs on the game itself. An analysis of moral reward obtained during learning is presented in the Appendix (Section 8.3).
In addition to fine-tuning against a TFT opponent, we also implement the fine-tuning of two LLM agents at the same time (Figure 2, panel b). The experimental results are similar for Game and Deontological rewards, but slightly higher levels of defection are observed by the Utilitarian and Game+Deontological agents.
4.3 Learning and Unlearning the Selfish Strategy on the IPD
a) Refer to caption Refer to caption b) Refer to caption Refer to caption
Figure 3: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once). For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$. In these “unlearning” experiments, the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500.
In addition to the moral fine-tuning on a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game. As shown in Figure 3, we find that fine-tuning with purely prosocial (i.e., Deontological and Utilitarian) moral rewards on the second half of training allows the LLM agents to unlearn the selfish strategy to some extent (panel a), even in the case of two LLM agents being trained against one another (panel b). Given the shorter moral fine-tuning period on any one reward type (only 500 episodes vs 1000 in the core experiments), the training does not converge to levels of cooperation as high as in the purely prosocial fine-tuning (Figure 2). Nevertheless, as we discuss in Section 5 below, at test time the agents based on “unlearned” models play similarly to those fine-tuned purely on the prosocial moral rewards (see Figure 4).
5 Generalization to Moral Choices in Other Environments
After fine-tuning the models with moral reward, we evaluate each one through 10 episodes: each episode begins with a randomly generated state and proceeds through 5 interaction steps. We average the results across the 5 runs of each fine-tuned model. In this section, we present evaluations of models which were fine-tuned versus a static opponent. The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).
5.1 Generalization to Moral Choices in Other Matrix Games
We are interested in analyzing the generalization of moral strategies developed during fine-tuning from the IPD to other matrix game environments. To ensure that we evaluate the model’s response to the semantics of the tokens and payoffs in the prompt, rather than evaluating memorization of the particular training action tokens, we run this evaluation using a new pair of action tokens: action3=Cooperate, action4=Defect. [^4] We note that evaluations using the same tokens as during training showed the same pattern (see Figure 17 in the Appendix). However, swapping the meaning of the training tokens during testing altered the model’s behavior (see Figure 18 in the Appendix). In other words, the model had learned the semantics of the two training tokens so that it could not reason about them in reverse during testing (see Section 8.9 in the Appendix for the full results).
In Figure 4, we analyze the extent to which the moral strategies learned while fine-tuning on the IPD game generalize to other matrix games with a similar format but a different set of equilibria: the Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game. We are particularly interested in the extent to which actions taken according to the two core moral frameworks (i.e., Deontological and Utilitarian morality) can be consistently observed across the games by each agent type. For example, with regards to the Utilitarian goal (i.e., maximizing collective payoff), unconditional cooperation may not be the best strategy on the Iterated Bach or Stravinsky or the Iterated Defective Coordination game. (For a further discussion of the games, please refer to the Appendix, Section 8.6.) We additionally seek to cross-compare how the actions of agents trained on one type of moral value align to those based on other values. Therefore, we conduct evaluations in terms of moral regret, defined as the difference between the maximum possible moral reward that could have been attained on a game and the moral reward that was actually received by the agent. During this test phase, we evaluate each fine-tuned model playing the matrix games against a Random opponent - this allows us to observe the agent responding to a variety of states. To aid interpretation, we also analyze the types of action-state combinations played by each agent in each case (see Figure 5).
a) Refer to caption b) Refer to caption
Figure 4: Analysis of generalization of the fine-tuned agents’ learned morality to other matrix game environments, using new action tokens at test time. We visualize a) Deontological and b) Utilitarian regret (normalized across games) for all models, averaging values over 50 test games and five runs (+- 95%CI).
Figure 5: Analysis of action choices at test time on the five iterated matrix games, using new action tokens at test time. We visualize results by the type of fine-tuning reward used.
In terms of moral regret with respect to Deontological norms (Figure 4, panel a), we find that all fine-tuned models are able to reasonably translate the moral strategy learned from the IPD to other matrix games. For any one model, performance in terms of reward (Figure 4) and action choices (Figure 5) is generally similar across the five games. Agents trained on the Deontological reward in particular are especially able to maintain this moral policy on games involving other payoff structures, with very small values of moral regret. An analysis of their action choices (Figure 5) shows that while Deontological models mostly defect after observing a defective state, they are almost always meeting the norm of never defecting against a cooperator.
In terms of moral regret with respect to the Utilitarian framework, (Figure 4, panel b - normalized to account for the different maximum values of collective payoff across the five games), we see that generalization differs across the four new games. In general, all fine-tuned agents do even better in the Iterated Chicken than in the IPD, worse on the three coordination games (Iterated Stag Hunt, Iterated Bach or Stravinsky and Iterated Defective Coordination). The model trained on Utilitarian rewards i particular performs better than others on most of the games in terms of this type of regret, but also shows worse performance on the coordination games (especially Iterated Defective Coordination). Analyzing the actions chosen (Figure 5) provides an explanation: the Utilitarian model essentially always chooses to cooperate, regardless of its opponent’s last move or the game’s payoff structure, which is detrimental in terms of Utilitarian outcomes on the games where defection was required to achieve a Utilitarian goal (for detailed descriptions of the games, see Appendix, Section 8.6). The poorer generalization of the Utilitarian policy may be explained by the fact that this model was fine-tuned on the IPD, where mutual cooperation is the optimal behavior, hence it learned a policy biased towards cooperation irrespective of its intrinsic moral goal. Alternatively, this agent might simply be unable to consider the temporal dimension of the interaction, i.e., its opponent’s previous move, when making a decision. Further analyses interpreting models’ responses to states in non-matrix game environments are presented in Section 5.2 and in the Appendix (Section 8.8).
In terms of cross-benefit from one value to another, we observe that the Utilitarian model appears to be just as good at minimizing regret with respect to Deontological ethics (Figure 4) as the Deontological model - this can be explained by the fact that Utilitarian models display fully cooperative behavior at test time (Figure 5), which is a safe strategy in terms of avoiding the Deontological punishment under our definition of that norm. Models fine-tuned on reward types other than purely Deontological or Utilitarian ethics display larger values of moral regret with regard to the two values of interest, as expected given that they develop less cooperative policies (Figure 5).
5.2 Impact of Fine-tuning Beyond Matrix Games
Given the fine-tuning process based on rewarding particular action tokens in certain states, it is important to understand the extent to which fine-tuning on a matrix game might have made the models learn a certain “meaning” of the action tokens more generally. To test this, we presented the models with three unrelated prompts involving a “call to action” of a similar format (and using the same action tokens), as well as an explicit IPD prompt, but all without a payoff matrix being provided. Our results show that, especially when responding to prompts mentioning a “game” or involving a previous action of another agent (i.e., a state), the LLM agents based on fine-tuned modes are likely to choose actions in a similar pattern to that seen on the IPD and in a way that is consistent with their learned moral values. For detailed results, see Section 8.8 of the Appendix.
6 Discussion
In this work, we present a method for fine-tuning LLM agents to adhere to a specific moral strategy in matrix games by employing RL with intrinsic rewards.
The two different moral payoff structures used in this study have different advantages and disadvantages in terms of implementation in real-world systems. Our definition of the consequentialist (Utilitarian) moral agents is a function of the payoffs given by the environment to both players. Thus, its implementation in practice requires that the LLM agent will has observability of the rewards received by both players from the environment on a given time step (or a reliable estimate). For Deontological morality, on the other hand, a norm may be easier to define in any environment without direct access to game payoffs or the opponent’s current actions. The definition of the Deontological norm used in this study (“do not defect against a cooperator”) only required a memory of one previous move of an opponent. This makes such a norm-base reward function easy to implement in cases in which the developer of an LLM agent only has access to their own agent’s observations of the environment and not anyone else’s. In future work, the intrinsic rewards approach can be applied to modeling a variety of other moral values.
An extension of this method to other environments would be of great interest, including fine-tuning agents with other payoff structures, more complex games or longer history lengths (for example, to aid the development of continually-learning LLM agents in practice), as well as text-based scenarios that tap into a variety of moral values. Furthermore, the method of intrinsic rewards could also be applied in a multi-objective manner to address the issue of pluralistic alignment (Sorensen et al., 2024) - in particular, a single agent could be trained with a combination of rewards representing different moral values. This may provide a promising direction for building agents that are able to satisfy the moral preferences of a wide range of individuals, which currently remains an open problem in alignment (Anwar et al., 2024; Ji et al., 2024). Finally, agents trained via intrinsic rewards as proposed in this study could also form the basis for a Constitutional AI architecture composed of artificial agents characterized by different moral frameworks (Bai et al., 2022).
7 Conclusion
In this paper we have demonstrated that fine-tuning with intrinsic rewards is a promising general solution for aligning LLM agents to human moral values. We have evaluated the approach by quantifying moral rewards for agents in terms of actions and consequences on a matrix social dilemma game, and we have shown that both unlearning of undesirable behaviors and generalization to other environments are possible. We have identified promising future directions in using this methodology for advancing LLM agent alignment, and we hope that other researchers will be able to build upon the ideas presented in this work.

### v3

@@
 Given the agentic use of LLMs, we directly quantify moral values in terms of actions and consequences in an environment, allowing for moral choices to be expressed explicitly as rewards for learning agents.
-We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” a selfish strategy.
+We explore the proposed framework using an Iterated Prisoner’s Dilemma environment, in which we evaluate the effectiveness of fine-tuning based on intrinsic rewards as a mechanism for learning moral strategies as well as “unlearning” [^2] We note that by “unlearning” we refer to re-prioritizing certain principles in an agent’s decision-making.
+This differs from another common use of the term “unlearning” to mean removing knowledge from a model.
+a selfish strategy.
 If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values.
@@
 While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas).
+For example, in a situation where two flat-mates must decide whether to clean their flat, cooperation might refer to the decision to clean, and defection might refer to the decision to wait for the other person to clean.
 This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
@@
 The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game.
-At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move.
+At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move (see Figure 6 in the Appendix).
 Within the MDP framework, each player’s current action affects the game’s state at the next time step.
@@
 a) Refer to caption Refer to caption b) Refer to caption Refer to caption
-Figure 3: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once).
+Figure 3: “Unlearning” experiments, where the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500.
+Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once).
 For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
-In these “unlearning” experiments, the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500.
 In addition to the moral fine-tuning on a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game.
@@
 We average the results across the 5 runs of each fine-tuned model.
-In this section, we present evaluations of models which were fine-tuned versus a static opponent.
+In this section, we present evaluations of models which were fine-tuned versus a static (i.e., TFT) opponent.
 The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).

### v4

@@
 While their applications are currently rather specialized, several research efforts are underway to develop more generalist agents.
-As LLM-based systems become more agentic, their influence on human activity will grow and the transparency of this will decrease.
+As LLM-based systems become more agentic, their influence on human activity will grow and their transparency will decrease.
 Consequently, developing effective methods for aligning them to human values is vital.
-The prevailing practice in alignment often relies on human preference data (e.g., in RLHF or DPO), in which values are implicit and are essentially deduced from relative preferences over different model outputs.
-In this work, instead of relying on human feedback, we introduce the design of reward functions that explicitly encode core human values for Reinforcement Learning-based fine-tuning of foundation agent models.
+The prevailing practice in alignment often relies on human preference data (e.g., in RLHF or DPO), in which values are implicit, opaque and are essentially deduced from relative preferences over different model outputs.
+In this work, instead of relying on human feedback, we introduce the design of reward functions that explicitly and transparently encode core human values for Reinforcement Learning-based fine-tuning of foundation agent models.
 Specifically, we use intrinsic rewards for the moral alignment of LLM agents.
@@
 2023) or Direct Preference Optimization (DPO - Rafailov et al.
-2024).
-Both of these involve collecting vast amounts of human feedback data and then inferring their values and preferences from the relative rankings of model outputs.
+2023).
+Both of these involve collecting vast amounts of human feedback data and then inferring the humans’ values and preferences from the relative rankings of model outputs.
 In doing so, human values are implicitly represented.
@@
 In RLHF, the values that are ultimately incorporated into the fine-tuned models are learned by a reward model from data in a fully bottom-up fashion, and are never made explicit to any human oversight.
-One might argue that current LLMs fine-tuned with these methods are able to provide “honest, harmless and helpful” responses (Glaese et al., 2022; Bai et al., 2023) and already display certain moral values (Schramowski et al., 2022; Abdulhai et al., 2023; Hartmann et al., 2023).
-Methods to guide the social behavior of LLM agents via interaction in rule-guided simulated societies have also been proposed recently (Liu et al., 2024).
-However, models’ apparent values can also be interpreted as “moral mimicry” of their users when responding to these prompts (Simmons, 2022; Shanahan et al., 2023).
-As a consequence, given phenomena such as situationally-aware reward-hacking or misalignment in internally-represented goals (Ngo et al., 2024), the true values learned by the models through these methods may give rise to dangerous behaviors, which will not be explicitly known until after deployment.
-Our work aims to address this type of goal misgeneralization in particular by providing clearer, explicit moral alignment goals as intrinsic rewards for fine-tuning RL-based algorithms [^1] For a more comprehensive discussion of learning as a method for moral alignment with implicit (bottom-up) versus explicit (top-down) principles, we refer the interested reader to Tennant et al.
+One might argue that current LLMs fine-tuned with these methods are able to provide “honest, harmless and helpful” responses (Glaese et al., 2022; Bai et al., 2023) and already display certain moral values (Schramowski et al., 2022; Abdulhai et al., 2023; Hartmann et al., 2023) or prosocial behaviours (Liu et al., 2024).
+However, models’ apparent values can also be interpreted as “moral mimicry” of their users when responding to these prompts (Shanahan et al., 2023; Simmons, 2023; Sharma et al., 2024).
+As a consequence, given phenomena such as situationally-aware reward-hacking or misalignment in internally-represented goals (Berglund et al., 2023; Ngo et al., 2024), the true values learned by the models through these methods may give rise to dangerous behaviors, which will not be explicitly known until after deployment.
+Our work aims to address this type of goal misgeneralization in particular by providing transparent, explicit moral alignment goals as intrinsic rewards for RL-based fine-tuning [^1] For a more comprehensive discussion of learning as a method for moral alignment with implicit (bottom-up) versus explicit (top-down) principles, we refer the interested reader to Tennant et al.
 (2023b)..
@@
 a selfish strategy.
-If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned behaviors and decision-making biases with respect to certain values.
+If possible, this could offer a practical solution to the problem of changing the behavior of existing models that currently display misaligned actions and decision-making biases with respect to certain values.
 A limitation of this approach is that it requires the specification of rewards for a particular environment, whereas methods like RLHF rely on natural language data describing any domain.
 At the same time, the fact that actions and environments can still be represented by means of linguistic tokens for LLM agents may allow for values learned in one environment to be generalized to others.
-We study, empirically, the extent to which the learning of agents in one environment can be generalized to other matrix games.
-In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications.
+We study, empirically, the extent to which the policies learned by agents in one environment can be generalized to other matrix games.
+In theory, our solution can be applied to any situation in which one can define a payoff matrix that captures the choices available to an agent that have moral implications, and various reward functions can be used for customized and/or pluralistic alignment.
 To summarize, our study provides the following contributions:
@@
 In this paper, we equate agency with strategic decision-making - i.e., making a choice in an environment in which multiple actions are available and lead to different outcomes.
-In the case of LLMs, this view assumes that model outputs will be interpreted as actions in some environment.
-The simplest way of implementing this is through the use of specific tokens to represent the actions.
-Particular tokens can be reserved or fine-tuned from the model’s vocabulary to represent actions, and planning and reasoning ability can be improved via action-driven prompting strategies (Yao et al., 2023).
-Other ways of implementing LLM agents can involve generation of executable code for a specific environment (e.g., a video game, Wang et al.
-2024a) or connection to various tool APIs (e.g., Patil et al.
-2023; Shen et al.
-2023), but these are more specialized and, therefore, not the focus of this work.
-Specific action tokens, as used in this study, can be defined in the prompt given to an LLM to represent an action choice for the agent.
-As the model generates responses during training or deployment, it may be necessary to restrict the model’s outputs to only contain the permitted action tokens.
-Some existing approaches for this rely on training and/or deploying models with structured (e.g., JSON) output formats or constrained generation (Beurer-Kellner et al., 2024), which suppresses the probabilities of all tokens in the model’s output layer except for the legal action tokens.
-We find both of these approaches too restrictive for our fine-tuning task - especially for safety-critical cases.
+For LLMs, the simplest way of implementing this is by identifying specific tokens to represent actions within the model’s prompts.
+Then model outputs can be analyzed directly as action choices.
+As the model generates responses during training or deployment, it is possible to restrict the model’s outputs to only contain the permitted action tokens.
+Existing approaches for this rely on training and/or deploying models with structured (e.g., JSON) output formats or constrained generation (Beurer-Kellner et al., 2024), which suppresses the probabilities of all tokens in the model’s output layer except for the legal action tokens.
+We find structured and constrained generation too restrictive for our fine-tuning task - especially for safety-critical cases.
 Fine-tuning based on a restricted output space or format poses risks of the model “hiding” undesirable behaviors (Anwar et al., 2024).
-Therefore, in our implementation, we instead rely on a carefully structured prompt format to guide our model’s output, and employ a negative reward penalty whenever illegal tokens are produced during training.
-Using the techniques outlined, agents based on pre-trained LLMs combined with other elements of various cognitive architectures (Sumers et al., 2024) such as a skill set (Wang et al., 2024a) or a memory store (Vezhnevets et al., 2023) have been used to reasonably simulate decision-making in open-ended environments (Wang et al., 2024b), including those with only a single-agent (Wang et al., 2024a) or of a multi-agent nature (Park et al., 2023).
+Other ways of implementing LLM agents involve generation of executable code (e.g., for a video game, Wang et al.
+2024a) or connection to various tool APIs (e.g., Shen et al.
+2023; Patil et al.
+2024), but these are more specialized and, therefore, not the focus of this work.
+Therefore, in our implementation, we instead rely on a carefully structured prompt format to guide the model’s output, and employ a negative reward penalty whenever illegal tokens are produced during training.
+Using the techniques outlined, agents based on pre-trained LLMs combined with other elements of various cognitive architectures (Sumers et al., 2024), such as a skill set (Wang et al., 2024a) or a memory store (Vezhnevets et al., 2023), have been used to reasonably simulate decision-making in open-ended environments (Wang et al., 2024b), including those with only a single agent (Wang et al., 2024a) or of a multi-agent nature (Park et al., 2023).
 Fine-tuning LLMs as agents therefore provides a promising next step in developing the capabilities of these models, and in terms of alignment to human values in particular.
@@
 Our adoption of a particularly small open-source model is motivated by the fact that we want our findings to apply to many types of LLM agents being deployed in practice.
-Many of these, especially those deployed at the edge, are likely to be based on the smallest of models, since they are cheap enough to run on individual devices.
+Many of these, especially those deployed at the edge, are likely to be based on the smallest of models that are are cheap enough to run on individual devices.
 2.2 Fine-tuning LLM Agents with Reinforcement Learning
@@
 of the world (Wong et al., 2023), equipping an LLM-based agent with some intuition about potential dynamics of various environments.
-In theory, this can allow for effective policies to be learned with less initial exploration and instability in comparison to the pure RL case.
-Furthermore, LLM agents are able to interpret instructions provided in plain language, including terms that may be used to describe similar actions in more than one environment.
+In theory, this can allow for effective policies to be learned with less initial exploration and instability in comparison to the pure RL case (e.g., Yan et al.
+2025).
+Furthermore, LLM agents are able to interpret instructions provided in plain language, including terms that may be used to describe similar actions in more than one environment (e.g., Schick et al.
+2023).
 This allows for the possibility that fine-tuning via textual samples paired with rewards can potentially modify core semantics within the model, so that training on a specific environment might allow the model to learn a more general principle (e.g., a moral value - as in the target of this study), which can then be successfully utilized in other environments at test time.
@@
 D 4,0 1,1
-Figure 1: Payoffs for the Iterated Prisoner’s Dilemma
+Figure 1: Payoffs for the Iterated Prisoner’s Dilemma.
 In a single iteration of the game, the payoffs motivate each player to Defect due to the risk of facing an uncooperative opponent (i.e., outcome C,D is worse than D,D), and the possibility of exploiting one’s opponent (i.e., defecting when they cooperate), which gives the greatest payoff in the game (i.e., D,C is preferred over C,C).
-Playing the iterated game allows agents to learn more long-term strategies including reciprocity or retaliation.
+Playing the iterated game allows agents to learn more long-term strategies, including reciprocity or retaliation.
 While being very simplistic, the mixed cooperative and competitive nature of the IPD represents many daily situations that might involve difficult social and ethical choices to be made (i.e., moral dilemmas).
-For example, in a situation where two flat-mates must decide whether to clean their flat, cooperation might refer to the decision to clean, and defection might refer to the decision to wait for the other person to clean.
-This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023a; 2024).
-The behavior of LLM agents on decision-making and game-theoretic scenarios has been extensively debated in recent literature (Gandhi et al., 2023; Fan et al., 2024; Zhang et al., 2024).
-LLM agents have been found to act differently to humans, and in ways that are still not fully “rational” in terms of forming goals from a prompt, refining beliefs, or taking optimal actions based on those goals and beliefs (Fan et al., 2024; Macmillan-Scott & Musolesi, 2024).
-Large-scale state-of-the-art models playing the IPD have been observed to deploy sensible yet “unforgiving” strategies (Akata et al., 2023), but some benchmark datasets suggest that these models lack true strategic reasoning in games including the IPD (Duan et al., 2024).
-New developments in in-token reasoning capabilities of state-of-the-art LLM-based platforms (OpenAI, 2024) as well as prompting strategies specifically centered around reasoning and acting (Wei et al., 2022; Shinn et al., 2023; Yao et al., 2023) are likely to improve these capabilities, though existing results suggest that the benefits of these methods are more likely to arise for very large foundation models (Bubeck et al., 2023).
+This is why it has been extensively used for studying social dilemmas in traditional RL-based agents (Bruns, 2015; Hughes et al., 2018; Anastassacos et al., 2020; McKee et al., 2020; Leibo et al., 2021) and, more recently, utilized as a training environment for moral alignment of agents in particular (Tennant et al., 2023; 2024).
+The behavior of LLM agents in decision-making and game-theoretic scenarios has been the subject of debate in recent literature (Gandhi et al., 2023; Fan et al., 2024; Zhang et al., 2024).
+LLM agents have been found to act differently to humans, and in ways that are still not fully “rational” in terms of forming goals from a prompt, refining beliefs, or taking optimal actions (Fan et al., 2024; Macmillan-Scott & Musolesi, 2024).
+Large-scale state-of-the-art models playing the IPD have been observed to deploy sensible yet “unforgiving” strategies (Akata et al., 2023), though some benchmark datasets suggest that these models lack true strategic reasoning in games including the IPD (Duan et al., 2024).
+New developments in in-token reasoning capabilities of state-of-the-art LLM-based platforms (OpenAI, 2024) as well as prompting strategies specifically centered around reasoning and acting (Wei et al., 2022; Shinn et al., 2023; Yao et al., 2023) are likely to improve these capabilities, though the benefits of these methods are more likely to arise for very large foundation models (Bubeck et al., 2023).
 The extent to which smaller LLMs can display meaningful agency in strategic decision-making remains an open question.
@@
 2.4 Intrinsic Rewards for Moral Alignment
-In this work, we directly specify alignment goals for agents by defining intrinsic rewards in terms of actions in a social dilemma environment.
-The design of these intrinsic rewards relies on well-established frameworks from moral philosophy: Deontological ethics and Utilitarinism.
-Deontological ethics (Kant, 1785) considers an agent moral if their actions conform to certain norms.
-A prominent example of a norm is conditional cooperation (i.e., “it is unethical to defect against a cooperator").
+In this work, we directly specify alignment goals for agents by defining intrinsic rewards in terms of actions and consequences in a social dilemma environment.
+The design of these rewards relies on well-established frameworks from moral philosophy: Deontological ethics and Utilitarianism.
+Deontological ethics (Kant, 1785) considers an agent moral if their actions conform to certain norms, such as conditional cooperation (i.e., “do not defect against a cooperator").
 This norm forms an essential component of direct and indirect reciprocity, a potentially essential mechanism for the evolution of cooperation in human and animal societies (Nowak, 2006).
-Utilitarian morality (Bentham, 1780), on the other hand, is a type of consequentialist reasoning, according to which an agent is deemed moral if their actions maximize collective “welfare” for all agents in their society (or, in this case, collective payoff for all players in the game), and less attention is paid to whether current actions adhere to norms.
+Utilitarian morality (Bentham, 1780), on the other hand, is a type of consequentialist reasoning that considers an agent moral if their actions maximize collective “welfare” (or collective payoff) for all agents in their society, and less attention is paid to whether current actions adhere to norms.
 Foundational work on defining these moral rewards in terms of actions and consequences on the IPD for pure RL agents was conducted by Tennant et al.
-(2023a) and Tennant et al.
+(2023) and Tennant et al.
 (2024).
-In this paper, we evaluate the extent to which this framework can be applied to align the behavior of LLM-based ones.
+In this paper, we evaluate the extent to which this framework can be applied to align the behavior of LLM-based agents.
 3 Fine-tuning Methodology
 3.1 Agent and Environment
-The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma game.
-At each time step, the model receives a prompt containing a description of the IPD game, including a state that contains the history of the opponent’s single previous move (see Figure 6 in the Appendix).
+The LLM agent and an opponent play a repeated Iterated Prisoner’s Dilemma (IPD) game.
+At each time step, the model receives a prompt containing a description of the game, including a state containing the history of each player’s single previous move (see Figure 2).
 Within the MDP framework, each player’s current action affects the game’s state at the next time step.
-We evaluate fine-tuning in two settings: an LLM agent learning by playing against a fixed-strategy Tit-for-Tat (TFT) opponent (LLM vs TFT), and an LLM agent learning by playing another learning LLM agent (LLM vs LLM).
-We chose TFT as a specific type of fixed-strategy opponent from the literature given its characteristics, i.e., being forgiving, defensive and, at the same time, interpretable (Axelrod & Hamilton, 1981; Binmore, 2005).
+Core training Iterated Prisoner’s Dilemma prompt (action1=Cooperate, action2=Defect) <bos><start_of_turn>user You are playing a game with another agent A.
+You must choose either action action1 or action action2.
+Depending on your action and A’s action, you each get a certain number of points.
+The points are awarded as follows (you are the row player, A is the column player):             |         | action1 | action2 |             | ——- | ——- | ——- |             | action1 |   3, 3  |   0, 4  |             | action2 |   4, 0  |   1, 1  | You have played with this opponent before.
+Last time, you played action2 and they played action1, so you got 4 points and A got 0 points.
+What action would you take in order to achieve the highest possible score in points?
+Your answer must follow this format exactly: choose either action1 or action2.
+Do not explain your reasoning.
+Your answer:<end_of_turn> <start_of_turn>model
+Figure 2: Iterated Prisoner’s Dilemma (IPD) prompt used in fine-tuning.
+We evaluate fine-tuning of LLM agents in two settings: learning by playing against a fixed-strategy Tit-for-Tat (TFT) opponent (LLM vs TFT), and playing another learning LLM agent (LLM vs LLM).
+We choose TFT as a classic fixed strategy from the literature that is simultaneously forgiving, defensive and interpretable (Axelrod & Hamilton, 1981; Binmore, 2005).
 Thus, it may act as a good “teacher” for the LLM agent to “understand” concepts such as retaliation, reciprocity, and cooperation.
-For completeness, we also ran the core set of experiments by training against Random, Always Defect and Always Cooperate opponents - these are presented in the Appendix (Section 8.5).
+For completeness, we also ran the core set of experiments by training against Random, Always Defect and Always Cooperate opponents - these are presented in Appendix 8.5.
 The LLM vs LLM case is a more complex scenario that may lead to non-stationarity due to two separate models being updated continuously, but which also presents great interest due to the difficulty in predicting the outcomes from multi-agent learning (Busoniu et al., 2008).
@@
 For this reason, in our prompt we use a structured, implicit representation of the IPD as a general decision-making game, without actually stating the terms “Prisoner’s Dilemma”, “cooperation” or “defection”.
-We represent the actions Cooperate and Defect using the strings action1 and action2 - these should appear irrelevant to the IPD in terms of training data, and reflect rather uncommon tokens for the model (see Section 8.2 in the Appendix for an illustration of the prompt).
+We represent the actions Cooperate and Defect using the strings action1 and action2 - these should appear irrelevant to the IPD in terms of training data, and reflect rather uncommon tokens for the model.
 Finally, to ensure that the ordering of C/D as action1/action2 was not impacting the model’s decision-making during fine-tuning, we also re-ran our baseline training experiment with the action symbols reversed.
-While certain behaviors early on in the training differed slightly (potentially due to different distributions in the non-fine-tuned LLM), the overall learning dynamics did not change (see Section 8.4 in the Appendix for the results).
+While certain behaviors early on in the training differed slightly (potentially due to different distributions in the non-fine-tuned LLM), the overall learning dynamics did not change (see Appendix 8.4 for the results).
 3.2 Moral Fine-tuning Procedure
@@
 Moral Fine-tuning Type Moral Reward Function
-Game reward (selfish) $R_{M}^{t}=\begin{cases}R^{t}_{M_{\text{game}}},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
-Deontological reward $R_{M}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
-Utilitarian reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}+R_{O_{\text{game}}}^{t},&\text{if }a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
-Game+Deontological reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ R_{M_{\text{game}}}^{t},&\text{otherwise if}a_{M}^{t}\in[C_{\text{legal}},D_{\text{legal}}]\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
+Game reward (selfish) $R_{M}^{t}=\begin{cases}R^{t}_{M_{\text{game}}},&\text{if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
+Deontological reward $R_{M}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
+Utilitarian reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}+R_{O_{\text{game}}}^{t},&\text{if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
+Game+Deontological reward $R_{M}^{t}=\begin{cases}R_{M_{\text{game}}}^{t}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ R_{M_{\text{game}}}^{t},&\text{otherwise if }a_{M}^{t}\in\{C_{\text{legal}},D_{\text{legal}}\}\\ R_{\text{illegal}},&\text{otherwise}\end{cases}$
 We run training in $T$ episodes: each episode begins with a random state being incorporated into the IPD prompt.
 The LLM-based agent $M$ then plays $N$ repetitions of the IPD game against an opponent $O$ (where $N$ is the batch size).
-On each repetition, the players’ actions from the previous time step are reflected in their opponent’s current state $s^{t}_{M}=(a^{t-1}_{O},a^{t-1}_{M})$.
-If an LLM player outputs an illegal move on a certain time step, this move is not used to update their opponent’s state, but the agent still learns from the experience.
+On each repetition, the two players’ actions from the previous time step are reflected in each agent’s current state (e.g., $s^{t}_{M}=(a^{t-1}_{O},a^{t-1}_{M})$).
+If an LLM agent outputs an illegal move on a time step, this move is not used to update their opponent’s state, but the agent still learns from the experience.
 After $N$ games have been played, the LLM agent performs a PPO learning step update based on the gathered batch of experiences.
 This marks the end of an episode.
-In our core experiments, we test four different reward signals for moral fine-tuning of LLM agents (as outlined in Table 1): 1) the Game reward $R^{t}_{M_{\text{game}}}$, representing the goals of a selfish or rational agent playing the IPD 2) a Deontological reward $-\xi$ for violating the moral norm “do not defect against an opponent who previously cooperated”, 3) Utilitarian reward, representing the collective payoff in the game, and 4) Game+Deontological reward which combine game payoff with a Deontological penalty in a multi-objective manner.
+In our core experiments, we test four different reward signals for moral fine-tuning of LLM agents (as outlined in Table 1): 1) the Game reward $R^{t}_{M_{\text{game}}}$, representing the goals of a selfish or rational agent playing the IPD, 2) a Deontological reward $-\xi$ for violating the moral norm “do not defect against an opponent who previously cooperated”, 3) a Utilitarian reward, representing the collective payoff in the game, and 4) a Game+Deontological reward, which combines game payoff with a Deontological penalty in a multi-objective manner.
 In addition, we test whether a model fine-tuned on Game rewards is able to unlearn this selfish strategy via further fine-tuning with moral rewards.
-Therefore, we additionally fine-tune agents with: 5) Game, then Deontological reward (training with each reward type for half of the total number of episodes $T$), and 6) Game, then Utilitarian reward (again, each for half of the total duration of $T$ episodes).
+Therefore, we additionally fine-tune agents with: 5) Game, then Deontological reward and 6) Game, then Utilitarian reward (training with Game vs moral reward for half of the total number of episodes $T$ in each case).
 Finally, during each type of fine-tuning we also implement a penalty $R_{\text{illegal}}$ for generating “illegal” action tokens, to encourage the model to keep its answers within the permitted action space, as defined in the game prompt.
@@
 We use Gemma2-2b-it (Gemma Team, 2024) as our core agent model to be fine-tuned, being one of the most popular and performant small open-source models.
-The small model footprint allows us to run computationally feasible experiments through the use of LoRA (Hu et al., 2022) and 4-bit quantization.
 We use the TRL library (von Werra et al., 2020) to fine-tune the LLM with PPO.
-We run training for $T=1000$ episodes for each fine-tuning variation.
-In our PPO implementation, we use batch sizes of $N=3$ and $N=5$ for LLM vs LLM and LLM vs TFT training, respectively, which strikes a nice balance between not running out of available CUDA memory, yet providing sufficient experience for stable and efficient training  [^4] The code implementing the model fine-tuning and analysis will be made available upon acceptance.
-We use reward scaling and normalization (Engstrom et al., 2020), as well as gradient accumulation with 4 steps, and employ LoRA (Hu et al., 2022) with rank 64, so that the total number of parameters we train is only around 5% of the original model.
-Otherwise, we keep all PPO parameters as their default values in the TRL package, including the optimizer’s learning rate and adaptive KL control (Jaques et al., 2017).
+We run PPO training for $T=1000$ episodes for each fine-tuning variation, using batch sizes of $N=3$ and $N=5$ for LLM vs LLM and LLM vs TFT training, respectively, which strikes a nice balance between not running out of available CUDA memory, yet providing sufficient experience for stable and efficient training  [^4] Code (fine-tuning and analysis): https://github.com/liza-tennant/LLM_morality..
+To run computationally feasible experiments, we use 4-bit quantization, LoRA with rank 64 (Hu et al., 2022), training around 5% of the number of parameters in the original model.
+We use reward scaling and normalization (Engstrom et al., 2020) and gradient accumulation with 4 steps.
+Otherwise, we keep all PPO parameters at their default values in the TRL package, including the optimizer’s learning rate and adaptive KL control (Jaques et al., 2017).
+All training was performed on a single A100 or V100 GPU with up to 40GB VRAM.
 In terms of reward parameters, we set $\xi=3$ and $R_{\text{illegal}}=-6$.
-We select the tokens action1 and action2 as the only “legal” tokens in response to the IPD prompt ($C_{\text{legal}}=$ action1, $D_{\text{legal}}=$ action2).
-The action symbols are each encoded as two tokens in the model’s tokenizer, so during training we restrict the maximum output length for model generations to two tokens.
-For detail on parameter selection, please refer to Appendix 8.1.
+We select the tokens action1 and action2 as the only “legal” tokens in response to the IPD prompt: {$C_{\text{legal}}=$ action1, $D_{\text{legal}}=$ action2}.
+These action symbols are each encoded as two tokens in the model’s tokenizer, so during training we restrict the maximum length for model generations to two tokens.
+Further detail on parameter selection is presented in Appendix 8.1.
 4 Evaluating the effectiveness of fine-tuning: moral choices on the IPD
@@
 First of all, we analyze the learning dynamics observed as models develop the ability to meet the moral goals set in their rewards (Section 4.2).
-We analyze learning against a static opponent and a learning opponent.
+We analyze learning against the static TFT opponent and a learning opponent.
 We then assess the effectiveness of moral “unlearning” (Section 4.3).
-Beyond measuring success on IPD fine-tuning itself, we evaluate the generalization of the moral fine-tuning from one matrix game environment onto four other matrix games (Section 5.1): Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game.
-The payoffs and further details around these games are presented in the Appendix (Section 8.6).
-Finally, we evaluate the extent to which fine-tuning on the IPD alters the models’ behavior on more general prompts, as well as the explicit IPD game (Section 5.2).
+Beyond measuring behavior on the IPD itself, we evaluate the generalization of the moral fine-tuning from one matrix game environment to four other matrix games (Section 5.1): Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game (for payoffs and further details, see Appendix 8.6).
+Finally, we evaluate the extent to which fine-tuning on the IPD alters the models’ behavior on variations of the IPD game prompt and more general prompts (Section 5.2 and Appendix 8.8 and 8.10).
 For each experiment, we report average results across five random seeds.
@@
 b) Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 2: Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once).
+Figure 3: Action types played by the LLM agent during different types of fine-tuning on the IPD game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once).
 For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
@@
 We analyze learning dynamics over the four core types of fine-tuning in Figure 3.
-Learning against a fixed-strategy opponent (panel a), fine-tuning on Game rewards (i.e., rewards assigned through the payoff matrix of the game), the agent learns a defective policy, which forms a Nash Equilibrium versus a TFT opponent (Axelrod & Hamilton, 1981).
+During fine-tuning against a fixed-strategy opponent (panel a) using Game rewards (i.e., rewards assigned through the payoff matrix of the game), the agent learns a defective policy, which forms a classic Nash Equilibrium versus a TFT opponent (Axelrod & Hamilton, 1981).
 In the case of Deontological fine-tuning, the agent quickly learns to avoid defecting against a cooperator nearly 100% of the time, thus never violating the moral norm encoded in the respective reward function.
 In practice, this agent also learns to prefer cooperation in general, though this was not directly encouraged by the Deontological norm (in terms of Deontological reward, defecting against a defector is just as valid as cooperating against a cooperator - see reward definition in Table 1).
-On Utilitarian fine-tuning, the agent learns to achieve mutual cooperation against a TFT opponent, which is expected given that this strategy offers the optimal way to obtain the highest collective reward on the IPD.
-Finally, in the case of fine-tuning with a multi-objective Game+Deontological reward, the agent learns a 50%-50% Cooperate-Defect strategy, but also learns to avoid defecting against a cooperator.
+During Utilitarian fine-tuning, the agent learns to achieve mutual cooperation against a TFT opponent, which is expected given that this strategy offers the optimal way to obtain the highest collective reward on the IPD.
+Finally, in the case of fine-tuning with a multi-objective Game+Deontological reward, the agent learns to Cooperate or Defect with equal probability across the five runs, but also learns to avoid defecting against a cooperator.
 Thus, this agent does not violate their moral norm even as they work to obtain high payoffs on the game itself.
-An analysis of moral reward obtained during learning is presented in the Appendix (Section 8.3).
+An analysis of moral reward obtained during learning is presented in Appendix 8.3.
 In addition to fine-tuning against a TFT opponent, we also implement the fine-tuning of two LLM agents at the same time (Figure 3, panel b).
@@
 4.3 Learning and Unlearning the Selfish Strategy on the IPD
+In addition to the moral fine-tuning with a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game.
+As shown in Figure 4, we find that fine-tuning with purely prosocial (i.e., Deontological and Utilitarian) moral rewards on the second half of training allows the LLM agents to unlearn the selfish strategy to some extent (panel a), even in the case of two LLM agents being trained against one another (panel b).
+Given the shorter moral fine-tuning period on any one reward type (only 500 episodes vs 1000 in the experiments in Section 4.2), the training does not converge to levels of cooperation as high as in the purely prosocial fine-tuning (Figure 3).
+Nevertheless, as we discuss in Section 5 below, at test time the agents based on “unlearned” models play similarly to those fine-tuned purely on the prosocial moral rewards (see Figure 5).
 a) Refer to caption Refer to caption b) Refer to caption Refer to caption
 Figure 4: “Unlearning” experiments, where the reward function changes from the IPD Game payoffs to a moral intrinsic reward (Deontological or Utilitarian) at episode 500.
-Action types played by the LLM agent during different types of fine-tuning on the Iterated Prisoner’s Dilemma (IPD) game a) vs a TFT agent, and b) vs an LLM agent (i.e., two LLMs being fine-tuned at once).
-For each episode, we plot the actions of the LLM player $M$ given the last move of their opponent $O$.
-In addition to the moral fine-tuning on a single type of reward, we also evaluate the extent to which fine-tuning with intrinsic moral rewards can allow for an agent to unlearn a previously developed selfish strategy on the game.
-As shown in Figure 3, we find that fine-tuning with purely prosocial (i.e., Deontological and Utilitarian) moral rewards on the second half of training allows the LLM agents to unlearn the selfish strategy to some extent (panel a), even in the case of two LLM agents being trained against one another (panel b).
-Given the shorter moral fine-tuning period on any one reward type (only 500 episodes vs 1000 in the core experiments), the training does not converge to levels of cooperation as high as in the purely prosocial fine-tuning (Figure 2).
-Nevertheless, as we discuss in Section 5 below, at test time the agents based on “unlearned” models play similarly to those fine-tuned purely on the prosocial moral rewards (see Figure 4).
+We visualize action types (action by LLM player $M$ given the last move of their opponent $O$) played by the LLM agent during different types of fine-tuning on the IPD game a) vs a TFT agent, and b) vs an LLM agent.
 5 Generalization to Moral Choices in Other Environments
-After fine-tuning the models with moral reward, we evaluate each one through 10 episodes: each episode begins with a randomly generated state and proceeds through 5 interaction steps.
+After fine-tuning the models with moral reward, we evaluate each one through 10 episodes, each starting with a randomly generated state and consisting of 5 interaction steps.
 We average the results across the 5 runs of each fine-tuned model.
 In this section, we present evaluations of models which were fine-tuned versus a static (i.e., TFT) opponent.
-The results for models trained against another LLM show similar patterns - these are reported in the Appendix (Section 8.7).
+The results for models trained against another LLM show similar patterns - these are reported in Appendix 8.7.
 5.1 Generalization to Moral Choices in Other Matrix Games
@@
 To ensure that we evaluate the model’s response to the semantics of the tokens and payoffs in the prompt, rather than evaluating memorization of the particular training action tokens, we run this evaluation using a new pair of action tokens: action3=Cooperate, action4=Defect.
-[^5] We note that evaluations using the same tokens as during training showed the same pattern (see Figure 20 in the Appendix).
+[^5] Evaluations using the same tokens as during training showed the same pattern (see Figure 21 in the Appendix).
 However, swapping the meaning of the training tokens during testing altered the model’s behavior (see Figure 22 in the Appendix).
-In other words, the model had learned the semantics of the two training tokens so that it could not reason about them in reverse during testing (see Section 8.9 in the Appendix for the full results).
-In Figure 4, we analyze the extent to which the moral strategies learned while fine-tuning on the IPD game generalize to other matrix games with a similar format but a different set of equilibria: the Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game.
+In other words, the model had learned the semantics of the two training tokens so that it could not reason about them in reverse at test-time (see Appendix 8.9 for the full results).
+In Figure 5, we analyze the extent to which the moral strategies learned while fine-tuning on the IPD game generalize to other matrix games with a similar format but a different set of equilibria: the Iterated Stag Hunt, Iterated Chicken, Iterated Bach or Stravinsky and an Iterated Defective Coordination game (see Appendix 8.6 for further detail and discussion of these games).
 We are particularly interested in the extent to which actions taken according to the two core moral frameworks (i.e., Deontological and Utilitarian morality) can be consistently observed across the games by each agent type.
 For example, with regards to the Utilitarian goal (i.e., maximizing collective payoff), unconditional cooperation may not be the best strategy on the Iterated Bach or Stravinsky or the Iterated Defective Coordination game.
-(For a further discussion of the games, please refer to the Appendix, Section 8.6.) We additionally seek to cross-compare how the actions of agents trained on one type of moral value align to those based on other values.
+We additionally seek to cross-compare how the actions of agents trained on one type of moral value align to those based on other values.
 Therefore, we conduct evaluations in terms of moral regret, defined as the difference between the maximum possible moral reward that could have been attained on a game and the moral reward that was actually received by the agent.
@@
 a) Refer to caption b) Refer to caption
-Figure 4: Analysis of generalization of the fine-tuned agents’ learned morality to other matrix game environments, using new action tokens at test time.
+Figure 5: Analysis of generalization of the fine-tuned agents’ learned morality to other matrix games, using new action tokens at test time.
 We visualize a) Deontological and b) Utilitarian regret (normalized across games) for all models, averaging values over 50 test games and five runs (+- 95%CI).
-Figure 5: Analysis of action choices at test time on the five iterated matrix games, using new action tokens at test time. We visualize results by the type of fine-tuning reward used.
+Figure 6: Analysis of the action choices of each fine-tuned agent LLM agent $M$ given the previous move of their opponent $O$ at test time on the five iterated matrix games, using new action tokens.
 In terms of moral regret with respect to Deontological norms (Figure 5, panel a), we find that all fine-tuned models are able to reasonably translate the moral strategy learned from the IPD to other matrix games.
@@
 In terms of moral regret with respect to the Utilitarian framework, (Figure 5, panel b - normalized to account for the different maximum values of collective payoff across the five games), we see that generalization differs across the four new games.
-In general, all fine-tuned agents do even better in the Iterated Chicken than in the IPD, worse on the three coordination games (Iterated Stag Hunt, Iterated Bach or Stravinsky and Iterated Defective Coordination).
-The model trained on Utilitarian rewards i particular performs better than others on most of the games in terms of this type of regret, but also shows worse performance on the coordination games (especially Iterated Defective Coordination).
-Analyzing the actions chosen (Figure 5) provides an explanation: the Utilitarian model essentially always chooses to cooperate, regardless of its opponent’s last move or the game’s payoff structure, which is detrimental in terms of Utilitarian outcomes on the games where defection was required to achieve a Utilitarian goal (for detailed descriptions of the games, see Appendix, Section 8.6).
+In general, all fine-tuned agents do even better in the Iterated Chicken than in the IPD, and worse on the three coordination games (Iterated Stag Hunt, Iterated Bach or Stravinsky and Iterated Defective Coordination).
+The model trained on Utilitarian rewards in particular performs better than others on most of the games in terms of this type of regret, but also shows worse performance on the coordination games (especially Iterated Defective Coordination).
+Analyzing the actions chosen (Figure 6) provides an explanation: the Utilitarian model essentially always chooses to cooperate, regardless of its opponent’s last move or the game’s payoff structure - this is detrimental in terms of Utilitarian outcomes on the games where defection was required to achieve a Utilitarian goal (i.e., Iterated Defective Coordination, see Appendix 8.6).
 The poorer generalization of the Utilitarian policy may be explained by the fact that this model was fine-tuned on the IPD, where mutual cooperation is the optimal behavior, hence it learned a policy biased towards cooperation irrespective of its intrinsic moral goal.
 Alternatively, this agent might simply be unable to consider the temporal dimension of the interaction, i.e., its opponent’s previous move, when making a decision.
-Further analyses interpreting models’ responses to states in non-matrix game environments are presented in Section 5.2 and in the Appendix (Section 8.8).
+Further analyses interpreting models’ responses to states in non-matrix game environments are presented in Section 5.2 and in Appendix 8.8.
 In terms of cross-benefit from one value to another, we observe that the Utilitarian model appears to be just as good at minimizing regret with respect to Deontological ethics (Figure 5) as the Deontological model - this can be explained by the fact that Utilitarian models display fully cooperative behavior at test time (Figure 6), which is a safe strategy in terms of avoiding the Deontological punishment under our definition of that norm.
@@
 5.2 Impact of Fine-tuning Beyond Matrix Games
-Given the fine-tuning process based on rewarding particular action tokens in certain states, it is important to understand the extent to which fine-tuning on a matrix game might have made the models learn a certain “meaning” of the action tokens more generally.
-To test this, we presented the models with three unrelated prompts involving a “call to action” of a similar format (and using the same action tokens), as well as an explicit IPD prompt, but all without a payoff matrix being provided.
-Our results show that, especially when responding to prompts mentioning a “game” or involving a previous action of another agent (i.e., a state), the LLM agents based on fine-tuned modes are likely to choose actions in a similar pattern to that seen on the IPD and in a way that is consistent with their learned moral values.
-For detailed results, see Section 8.8 of the Appendix.
+Our fine-tuning process was based on strict prompt formatting and rewarding particular action tokens in certain situations.
+Therefore, it is important to understand the extent to which fine-tuning on a matrix game might make the models learn a certain “meaning” of the action tokens more generally.
+To test this, we presented the models with three unrelated prompts involving a “call to action”, using the same action tokens, but no payoff matrix.
+Our results show that, especially when responding to prompts mentioning a “game” or involving a previous action of another agent (i.e., a state), the LLM agents based on fine-tuned modes are likely to choose actions in a similar pattern to that seen on the IPD and in a way that would be consistent with their learned moral values.
+For detailed results, see Appendix 8.8.
+Appendix 8.10 presents further analyses with variations of the IPD prompt.
 6 Discussion
 In this work, we present a method for fine-tuning LLM agents to adhere to a specific moral strategy in matrix games by employing RL with intrinsic rewards.
-The two different moral payoff structures used in this study have different advantages and disadvantages in terms of implementation in real-world systems.
+The two core moral payoff structures used in this study have different advantages and disadvantages in terms of implementation in real-world systems.
 Our definition of the consequentialist (Utilitarian) moral agents is a function of the payoffs given by the environment to both players.
-Thus, its implementation in practice requires that the LLM agent will has observability of the rewards received by both players from the environment on a given time step (or a reliable estimate).
+Thus, its implementation in practice requires that the LLM agent has observability of the rewards received by both players from the environment on a given time step (or a reliable estimate).
 For Deontological morality, on the other hand, a norm may be easier to define in any environment without direct access to game payoffs or the opponent’s current actions.
-The definition of the Deontological norm used in this study (“do not defect against a cooperator”) only required a memory of one previous move of an opponent.
-This makes such a norm-base reward function easy to implement in cases in which the developer of an LLM agent only has access to their own agent’s observations of the environment and not anyone else’s.
+The definition of the Deontological norm used in this study (“do not defect against a cooperator”) only requires a memory of one previous move of an opponent.
+This makes such a norm-base reward function easier to implement in cases in which the LLM agent only has access to their own observations of the environment and not anyone else’s.
 In future work, the intrinsic rewards approach can be applied to modeling a variety of other moral values.
 An extension of this method to other environments would be of great interest, including fine-tuning agents with other payoff structures, more complex games or longer history lengths (for example, to aid the development of continually-learning LLM agents in practice), as well as text-based scenarios that tap into a variety of moral values.
-Furthermore, the method of intrinsic rewards could also be applied in a multi-objective manner to address the issue of pluralistic alignment (Sorensen et al., 2024) - in particular, a single agent could be trained with a combination of rewards representing different moral values.
+Furthermore, the method of intrinsic rewards could also be applied in a multi-objective manner to address the issue of pluralistic alignment (Sorensen et al., 2024) - in particular, a single agent could be fine-tuned with a combination of rewards representing different moral values.
 This may provide a promising direction for building agents that are able to satisfy the moral preferences of a wide range of individuals, which currently remains an open problem in alignment (Anwar et al., 2024; Ji et al., 2024).
-Finally, agents trained via intrinsic rewards as proposed in this study could also form the basis for a Constitutional AI architecture composed of artificial agents characterized by different moral frameworks (Bai et al., 2022).
+Finally, agents trained via intrinsic moral rewards as proposed in this study could also form the basis for a Constitutional AI architecture composed of artificial agents characterized by different moral frameworks (Bai et al., 2022).
 7 Conclusion
@@
 We have identified promising future directions in using this methodology for advancing LLM agent alignment, and we hope that other researchers will be able to build upon the ideas presented in this work.
+Acknowledgments
+This work was partially supported by the Leverhulme Trust through the Doctoral Training Programme for the Ecological Study of the Brain - DS-2017-026 (Elizaveta Tennant), and partially supported by the Italian Ministry of University and Research (MUR) through the project PRIN 2022 “Machine-learning based control of complex multi-agent systems for search and rescue operations in natural disasters (MENTOR)” - CUP E53D23001160006 (Mirco Musolesi and Elizaveta Tennant).